OpenStudy (zyberg):

Prove that n! > n^2 for every integer n>= 4 and n! > n^3 for every integer n>= 6. Question is from induction chapter.

OpenStudy (zyberg):

In the example about Luca's sequence proof they used induction after dividing the number into the ones that it is made up of, however, I can't think of a way how to apply this technique to factorials...

OpenStudy (welshfella):

if n = k assume k! >= k^2 - which it is for k = 4 Now we need to show that when n = k+1:- (k+1)k! >= (k+1)^2

OpenStudy (welshfella):

Now can we derive that last one ?

OpenStudy (welshfella):

sorry that should be > not >=

OpenStudy (zyberg):

Hm... But how would it be possible to derive it? I understand the very basics, but I can't seem to understand how to get from k! > k^2 to (k+1)! > (k+1)^2 normally I would add something to both sides, but it's not a sequence, so I am having trouble...

OpenStudy (zyberg):

Oh, I see that it's possible to multiply both sides by (k+1), so the left side would become (k+1)! and right one - k^2(k+1), as you had said @welshfella. However, what would be the next step?

OpenStudy (welshfella):

what if we multiplied the first equation by (k + 1)

OpenStudy (welshfella):

lol1 - we both thought of that the same time!

OpenStudy (welshfella):

well since k is a positive number k^3 + k^2 is greater than (k + 1)^2

OpenStudy (faiqraees):

Base Case = n!>n² => H(k) Assuming H(k) is true n!>n² n! x n+1>n² x n+1 (n+1)!>n³+n²>(n+1)² for n>=4 H(k) => H(k+1) H(k) is true (prove that by plugging 4 into the base case)

OpenStudy (faiqraees):

And exactly the same strategy for the second part.

OpenStudy (welshfella):

Yes - thats basically the same reasoning.

OpenStudy (zyberg):

Oh! So, by proving that it is true for bigger number we proved that it is true for the lesser one... Interesting! Thank you all for help! :) (Not sure to whom should I give the medal now...)

OpenStudy (faiqraees):

Welshfella is the worthy candidate

OpenStudy (welshfella):

its up to you. I dont worry much about medals

OpenStudy (zyberg):

Okay, then Welshfella. Thank you again!

OpenStudy (welshfella):

yw